OpenStudy (anonymous):

You have $5 for lunch. You bought a turkey sandwich for $2.25. Write and solve and inequality to show how much more you can spend. Thank you!

OpenStudy (texaschic101):

2.25 + x < = 5 x < = 5 - 2.25 x < = 2.75

OpenStudy (anonymous):

Could you explain that to me?

OpenStudy (aveline):

Not exactly, remember to use the ≤ symbol @texaschic101

OpenStudy (aveline):

I can help explain

OpenStudy (anonymous):

Thanks for your time!

OpenStudy (aveline):

You have $5 to spend and have already spent $2.25. This can be shown by: x+2.25≤5 This means that whatever we spend (including the $2.25) must be equal to or less than $5. We can simplify the equation: x≤2.75 This means that we have x dollars left to spend, or $2.75.

OpenStudy (anonymous):

Thanks sooooo much! : )

OpenStudy (aveline):

No problem